OpenStudy (anonymous):

find the derivative,y=ln(cos^2theta) help please!

OpenStudy (campbell_st):

derivative if log(f(x)) is f'(x)/f(x) use the fact that \[\cos^2 \theta = (\cos(\theta))^2\] \[dy/dx = -2\sin(\theta)/(\cos^2 \theta)\] this can be simplified to \[dy/d \theta = -2\sin \theta/\cos \theta \times 1/\cos \theta\] or \[dy/d \theta = -2\tan \theta \sec \theta\]

OpenStudy (anonymous):

I think the derivative of cos^2 theta is:  so f'(x)/f(x) = -2*tan(theta)